### v1

# Quantifying task-relevant representational similarity using decision variable correlation

Abstract
Previous studies have compared the brain and deep neural networks trained on image classification. Intriguingly, while some suggest that their representations are highly similar, others argued the opposite. Here, we propose a new approach to characterize the similarity of the decision strategies of two observers (models or brains) using decision variable correlation (DVC). DVC quantifies the correlation between decoded decisions on individual samples in a classification task and thus can capture task-relevant information rather than general representational alignment. We evaluate this method using monkey V4/IT recordings and models trained on image classification tasks. We find that model–model similarity is comparable to monkey-monkey similarity, whereas model–monkey similarity is consistently lower and, surprisingly, decreases with increasing ImageNet-1k performance. While adversarial training enhances robustness, it does not improve model–monkey similarity in task-relevant dimensions; however, it markedly increases model–model similarity. Similarly, pre-training on larger datasets does not improve model–monkey similarity. These results suggest a fundamental divergence between the task-relevant representations in monkey V4/IT and those learned by models trained on image classification tasks.
1 Introduction
Deep learning [1, 2] has revolutionized how neuroscientists construct models of the brain. For vision science, deep neural networks offer appealing candidate models for studying the primate ventral pathway [3, 4, 5] and, more recently, dorsal pathway [6, 7, 8]. A decade ago, it was reported that the internal representations in convolutional neural networks (CNNs) trained on image categorization can explain a substantial fraction of variance in higher visual areas, surpassing the classic models for these brain areas [3]. Follow-up research has tested many more variants of deep networks on their aligment with the brain using both neural data [9, 10, 11] and behavior data [12, 13, 14] from monkeys and humans. Newer deep network models are learning to leverage information with increasing efficiency and reliability, mirroring the evolutionary journey of primate brains. Naturally, one appealing hypothesis is that deep networks that exhibit higher accuracy and robustness in vision tasks, or trained on larger datasets would better explain visual processing in the brain.
The key therefore lies in how to compare deep networks and the brain.
One class of methods seeks to quantify the similarity of internal representations between models and brains. This includes methods such as representational similarity analysis (RSA) analysis[15], linear regression[5, 16], and generalized shape metrics [17]. More recently, another class of methods that put more emphasize on quantifying the behavioral similarity has been proposed, including Cohen’s Kappa [13] and I2n behavioral predictivity [12, 18]. The goal of these methods is to provide a high-resolution, image-by-image comparison of the decision strategies used by neural networks and the brain. One challenge has been how to properly disentangle the model accuracy, decision biases and decision consistency from behavior [19, 13]. Interestingly, these methods seem to find contradictory trends in model-brain alignment. The reason for this remains unclear.
Here, we propose a principled approach that combines the merits of model comparisons at the representation and behavior levels. Our new approach specifically quantifies the trial-by-trial consistency of two neural representations for solving a classification task, ignoring features that are irrelevant for the task. In doing so, it enables us to isolate the consistency of the decision strategy of two observers when solving a behavioral task. Our method is based on decision variable correlation (DVC) developed to measure behavioral similarity based on choice data [19], and we have generalized it to analyze the consistency of high-dimensional neural representations. Applying our method to compare neural representations for solving image recognition tasks from monkey brains and deep network models led to several interesting findings. In particular, we found that model–model and monkey–monkey similarities are comparable, whereas model–monkey similarity is consistently lower and decreases with increasing ImageNet-1k accuracy. Somewhat surprisingly, this gap is not remedied by adversarial training or training on larger datasets.
2 Background and relevant work
Community efforts have pushed towards better methods to compare brains and models and for brain-model alignment. Different factors have been hypothesized to be relevant to the alignment, including model architecture, robustness, and training data, as summarized below.
Model architecture and scale One hypothesis has been that as models improve in task performance or architectural complexity, their internal representations become more brain-like. The Brain-Scores [18] of the image classification models were reported to be positively correlated to ImageNet-1k accuracy, although the trend plateaus at higher accuracy. On the other hand, studies using RSA reported that neither model scale nor architecture significantly improved alignment to human behavioral similarity judgments [9]. Another study using RSA reported negative correlation between alignment to human neural activity and model complexity (in FLOPs) [20]. Studies using Cohen’s Kappa reported that human-model behavioral consistency at the image level remains low despite improved performance on out-of-distribution datasets with scaling[13, 14].
Robustness In addition to good performance, the primate visual system is robust against external and internal noise, prompting the question of whether robustness to adversarial perturbations or corruptions is related to brain-model alignment [21]. By enforcing alignment with monkey IT representations, models exhibited both enhanced adversarial robustness and increased behavioral alignment with human subjects. Another study found that model metamers – artificial stimuli that elicit the same response as natural stimuli, generated by robust models– are more recognizable to humans, but is not itself predictive of recognizability [22]. However, studies using Cohen’s Kappa report that robust models still diverge from humans on their error patterns[14].
Rich and multimodal training data Using Cohen’s Kappa, [14] reports that models trained on larger and more diverse datasets become more human-like in their behaviors. On the other hand, a recent large-scale study using a variation of RSA reported that upgrading from ImageNet-1k to ImageNet21k does not significantly improve alignment to human brain, but object-oriented ImageNet datasets lead to much better alignment than datasets containing only places or faces [11]. Similarly, an ecologically-motivated dataset seems to improve model-brain alignment [23]. Joint vision-language models such as CLIP has also been shown to better predict human brain activity [14, 10, 24].
Similarity measures The reader might have already observed that different similarity measures between representations or behaviors can sometimes yield very different results. Indeed, recent studies warned that different methods could lead to different conclusions [25, 26]. Meanwhile, researches have been examining whether the classic similarity measures are indeed widely applicable: different models result in similar level of linear predicitivity of the brain [18, 11]; RSA is blind to the specific features used to solve a task [27]; later we also show that measures of behavioral similarity may be biased by the choice of decoders [13, 14].
3 DVC: Quantifying the trial-by-trial consistency of two representations
We develop a new method to evaluate the consistency of two representations. This method is based on a principled generalization of signal detection theory. It enables one to estimate how correlated the decision strategies of two observers are on a classification task. The method is insensitive to the observers’ biases and is not confounded by the behavioral accuracy. It operates at the level of neural representations, and enables one to analyze the internal representation to infer the consistency of the two representations for solving the classification tasks. Thus, the method can quantify task-relevant representational similarity. Compared to methods purely based on behavior [19, 13], it takes advantage of the richness of the internal representations of neural networks and brains. Meanwhile, in contrast to methods for analyzing the similarity of two neural representations (such as representational similarity analysis), our methods focus on the dimensions that are relevant for a behavioral task and is invariant to variability along other task-irrelevant dimensions.
Figure 1: The computational framework of decision variable correlation (DVC) for neural representations. (a) Traditional signal detection theory models how a single observer solve a binary classification task. The idea is that the observer use a decision variable together with a criterion (dash line) to make a choice. (b) Decision variable correlation generalizes the signal detection theory to study the trial-by-trial consistency of the decision variables of two observers. The two panels illustrate two cases with the same accuracy in solving the task, but with drastically different correlations in the decision variabless (DVs). (c) We further generalize DVC to compare two neural representations. The basic idea is to use optimal linear classifier to infer the decision variables of individual observers and then quantify the consistency of the decision variables.
3.1 Decision variable correlations (DVC) of two neural representations
Signal detection theory is fundamental in the study of perceptual behavior. The idea is that, for binary-choice tasks, observer uses a continuous decision variable (DV) to make a choice (Fig. 1a). Recently, [19] proposed to generalize signal detection theory to study the correlation of decision variables of two observers(Fig. 1b). Their method inferred the decision variable correlation from binary choice data. Here, we develop a simple new strategy to infer the trial-by-trial decision variable correlation from the high-dimensional internal representations (Fig. 1c).
For a pair of image categories and an observer (a brain area or a particular layer from a neural network), we can take its neural representation and find the optimal decision axis for solving the categorization task. We then project the high-dimensional representation for each image onto the decision axis and obtain its decision variable. Now consider the case of two observers. By performing the analysis on both observers, we obtain two decision variables for each image. We can compute the correlation of the decision variables (DVC) for the two observers (Fig. 1c). This correlation captures the similarity of the encoding and the decoding into a decision for the two observers in this classification task.
Note that the method of inferring DVC from behavioral responses only applies to binary choice tasks. Our new method does not suffer this limitation. Given N (>2) image classes, we can focus on each pair of categories at a time, and infer the DVC for that particular classification task.
3.2 Implementation of the method
We next discuss practically how we implement the DVC framework to analyze the high-dimensional neural representations from the brains or the deep networks.
Decoding decision variables (DVs) from neural representations For each pair of classes (e.g., cats v.s. dogs), we use Linear Discriminant Analysis (LDA) to find the axis that maximizes class separation to decode the DVs from the brain or model representations. The projection onto the LDA axis reflects the model’s tendency to classify the image as one class versus the other; values near the midpoint indicate greater classification uncertainty. One important practical question is that LDA can be unstable under high dimensions with few samples. The reason is that there are many noisy feature directions with similar class separation, but the projections of image representations along these dimensions can be different. Consequently, even if two models have the same underlying representations, LDA projections may show low correlation. Note that models examined in this paper have a wide range of dimensionality in their penultimate layer ($10^{3}-10^{7}$).
To address this problem, we use dimensionality reduction (e.g., PCA) to reduce the representations to the same number of features before using LDA to decode the underlying DV 11125 PC dimensions. See Appendix C for experiments that demonstrate the robustness of the results.. Importantly, the cross-validated LDA accuracies are high for all representations tested. We measure the similarity between decoded DVs using Pearson Correlation. A DVC value is obtained for each class in each pair of classes. The final reported number is taken as the average of all DVC values.
Normalization for correcting the measurement noise Different systems have different noise levels that limit the correlation that could be reliably decoded. The otherwise perfect correlation between two identical representations would be corrupted by added noise. Low correlation might therefore reflect true underlying dissimilarity or high noise level. To correct for the under-estimation of DVCs due to measurement noise, we develop a split-half procedure to infer the impact of noise.
We aim to estimate the true correlation between two decision variable (DV) signals, $\text{DV}_{A}$ and $\text{DV}_{B}$, each of which is contaminated by independent noise. To correct for the attenuation bias introduced by noise, we split each DV into two independent halves: $\text{DV}_{A1}$, $\text{DV}_{A2}$ and $\text{DV}_{B1}$, $\text{DV}_{B2}$. For neural recordings, this would indicate splitting into two sets of neurons, and for model representations, two sets of hidden units. We then compute a noise-corrected Pearson correlation as follows:
$\rho_{\text{corrected}}=\frac{r_{\text{cross}}}{r_{\text{self}}}$ (1)
where the numerator reflects the geometric mean of all pairwise cross-observer correlations:
$r_{\text{cross}}=\left[\rho(\text{DV}_{A1},\text{DV}_{B1})\cdot\rho(\text{DV}_{A1},\text{DV}_{B2})\cdot\rho(\text{DV}_{A2},\text{DV}_{B1})\cdot\rho(\text{DV}_{A2},\text{DV}_{B2})\right]^{1/4}$ (2)
and the denominator normalizes by the geometric mean of the within-observer (split-half) reliabilities:
$r_{\text{self}}=\left[\rho(\text{DV}_{A1},\text{DV}_{A2})\cdot\rho(\text{DV}_{B1},\text{DV}_{B2})\right]^{1/2}$ (3)
This correction removes the bias introduced by independent noise, yielding an unbiased estimate of the true underlying correlation between the latent signals driving $\text{DV}_{A}$ and $\text{DV}_{B}$. 222 The splits are generated randomly. See Appendix A which proves that this recovers the true underlying correlation. See Appendix C for discussions on behaviors of split normalization at boundary conditions.
4 DVCs reveal the divergence between deep networks and brains
We apply the new DVC-based methodology to examine (i) the trial-by-trial consistency of the neural representation of the macaque high-level visual areas (V4/IT), (ii) the consistency of the neural network models and the IT/V4 neural representations in macaque monkeys, as well as (iii) the consistency of different deep neural network models. We will specifically consider three classes of deep network models: (i) models that were pre-trained on ImageNet-1k using standard network training (i.e., no adversarial training); (ii) "robust models" that were fine-tuned on ImageNet-1k using adversarial training; (iii) "data-rich models" that were pre-trained on even larger datasets such as ImageNet-21k and JFT-300M.
Figure 2: Results on models trained on ImageNet-1k. (a) Heatmap: DVCs inferred for pairs of models. Different colors are used to indicate models from different model families. 15 models are selected to represent this cohort in later analysis. (b) 2D t-SNE embedding of the models using their dissimilarities, measured as $1-DVC$. (c) There is a strong negative correlation between the classification performance (top-1 accuracy) of a network and its DVC correlation to the V4/IT representation. (d) Networks belonging to the same family exhibit higher DVCs compared to those belonging to different model families (p = 1.33e-56).
4.1 High trial-by-trial consistency of V4 & IT representations across monkey brains
[…]
We combined the neural data from areas V4 and IT, and computed the DVCs. We find that the DVC between the monkeys is about 0.57. We further compute the DVCs for V4 and IT separately, and find the DVC values to be 0.63 and 0.41, respectively. Overall, these results suggest that DVCs across the monkeys’ brain are generally high, implying that the encoding and the decision strategies used by different monkeys are consistent on an image-by-image basis.
4.2 Deep networks with higher accuracy on ImageNet exhibit lower DVCs with brains
We study a set of models (n=43, obtained from Torchvision) [29] pretrained on ImageNet-1k, an influential benchmark in image classification. This also offers a fair comparison between models by controlling for confounding factors related to different training data. We used the same datsets from [28] as above. We feed the images in [28] into deep vision-based neural networks, subject to standard transforms. The model representation is defined as the activation in response to the image, taken from the penultimate layer – the last layer before the final logit layer.
[…]
Network vs. network We next examine the DVCs between different deep neural networks. Specifically, we evaluate DVCs between deep networks from different model families [^3] We define a model family as a set of architectures sharing a canonical computational graph – such as residual, attention, or convolutional block structures—with variation limited to hyperparameters like depth, width, patch size, or token embedding dimension. See Appendix B for more model details. Using DVC, models from the same family or otherwise share architectural similarities are clearly clustered together(Fig. 2a,b). We find that DVCs between models within the same family (similar model structures and training processes) are substantially higher than pairs from different families(p = 1.33e-56, Fig. 2d), aligning with previous findings [30]. We also find DVCs between models not to be exceedingly high. Despite being trained on the same dataset, they do not seem to converge to a single solution, at least not significantly higher compared to the similarity between the two monkeys (Fig. 2d). These results imply that model structures and training processes still play significant roles in the task solutions found by the models.
Notably, these results differ substantially from results obtained by computing error consistency[13, 14]. Geirhos et al. reported that (i) the consistency between model and brain is very low; (ii) the consistency between network models is much higher than the consistency between humans. Later, we will address the difference between the methodologies.
Figure 3: Results on robustly trained deep networks on ImageNet-1k. (a) Networks based on adversarial training has lower DVC with V4/IT compared to the representative models (introduced in Fig.2)) without adversarial training. (b) Heatmap showing the inferred DVCs between pairs of models. (c) Robustness networks have high DVCs among themselves, and they have relatively low DVCs with the representative models.
4.3 Adversarially trained networks, while highly consistent, have low DVCs to the brain
Robustness represents one important difference between deep networks and our perceptual systems. Small perturbations to images that are imperceptible to humans can lead to qualitative errors in deep networks (i.e., adversarial examples) [31]. Adversarial examples reflect the misalignment between representations in deep networks and brains, given certain local perturbations in the inputs. Adversarial robustness can be increased by using adversarial training, e.g., by finding adversarial examples and adding them to the training set. Studies suggest that features learned through adversarial training may be more aligned with human perception [32, 22], posing an intriguing hypothesis that by making networks locally consistent with human perception, network representations may be better aligned with brain representations globally.
[…]
These results suggest that adversarially trained models converge to a common solution (despite that these models have different architectures). Their representations diverge from the non-adversarially trained deep networks, and unfortunately, they also diverge from the representation in V4/IT.
Figure 4: Results on deep networks trained on richer datasets. (a) Networks we examined that were pre-trained on richer datasets exhibit lower DVC with V4/IT compared to the representative models (trained on ImageNet-1k). (b) Heatmap showing the inferred DVCs between pairs of models. (c) DVCs between lower data-rich models and representative models are generally lower than those within representative models or data-rich models.
4.4 Networks trained on rich datasets exhibit no increase in DVC to the brain
Whereas ImageNet-1k has been an important benchmark dataset for the image classification community for a decade, recent state-of-the-art models are trained on larger datasets such as ImageNet-21k, which is a scaled-up version of ImageNet-1k, and JFT-300M, which is proprietary. Models trained on larger, more diverse datasets may generalize within a larger domain, and may show better out-of-distribution generalization ability. A recent study showed that models trained on these larger datasets may exhibit better alignment with human behavior [14]. Furthermore, the negative correlations between classification performance and DVCs to the brains (Fig. 2c) suggest the possibility of overfitting to a particular dataset that is much smaller than what brains are trained on evolutionarily, developmentally and during the experiments[34]. Therefore, it is of particular interest to investigate whether models trained on the richer datasets exhibit higher DVCs to the brain.
[…]
Figure 5: Comparsion to Cohen’s Kappa. (a) Heatmaps showing the DVC and Cohen’s Kappa for pairs of representative models. (b) There is a strong positive correlation between Cohen’s Kappa and DVC on model-model consistency (evaluated on this dataset) (c) There is a decent positive correlation between Cohen’s Kappa and DVC on model-monkey consistency. (d) The response histogram to ’edges’ distortion based on the model and decision rules used in [13] and the original study [40]. Different colors represent different nerual networks. (e) Simulation results show that Cohen’s Kappa is sensitive to decision biases, while DVC is invariant to decision biases.
4.5 Comparison to error consistency based on Cohen’s Kappa
One method that is highly relevant to DVC is Cohen’s Kappa. As a classic statistical measure of inter-rater consistency [41], it was recently applied to quantify the error consistency of deep networks and brains [13, 14]. These studies arrived at very different conclusions, namely that model-model similarity is significantly higher than model-human similarity and that models trained on rich datasets are more aligned with humans. At a high level, these results seem to be inconsistent with our findings, because we found that (i) deep networks exhibit modest consistency with the brain; (ii) DVCs between different deep networks trained on ImageNet-1k are not exceedingly large; (iii) models trained on rich datasets have lower DVCs with brains.
[…]
What then is causing the difference in the network-network and network-brain consistences between the results reported in  [13] and DVC? According to signal detection theory, Cohen’s Kappa is determined by both the correlation of the underlying decision variables and the decision criterion. Thus, we wondered if the extremely high Cohen’s Kappa values between different networks as reported in  [13] were due the biases in the decisions. We thus performed a second analysis to further investigate the data from  [40] and analysis used in  [13]. Consistent with our hypothesis, we find that the approach in  [13] introduces high decision bias (see Fig.5d) and reduced accuracy, especially when target categories do not align cleanly with the original training labels. We also find that the origin of this large decision bias is because the analysis in [13] is based an aggregated decoder that estimates class probabilities by combining probabilities from related ImageNet-1k classes. Once we substituted the original decoder with a cross-validated logisitic regression classifier, the estimated Cohen’s Kappa values become largely consistent with the DVC we obtained on the main dataset we analyzed. Finally, using a simple simulation, we demonstrated that Cohen’s Kappa is inflated in this setting as bias increases and accuracy drops, whereas DVC remains consistent (Fig.5e).
These results also highlight a key advantage of the DVC method – it is insensitive to the decision biases. The error consistency quantified based on Cohen’s Kappa captures both shared behavior biases and consistency in their underlying decision variables. For future work, it would be interesting to combine the two methods to dissect the contribution of consistency of DVs and the shared biases in the observers. Doing so requires the presence of both neural and behavior data for the same set of stimuli.
5 Discussion
We have developed a new method, DVC, to quantify the consistency of the two neural representations that emphasizes task-relevant features. DVC is distinct from other popular approaches such as representational similarity analysis [15] or linear regression [5]. Two representations could have high DVC but low consistency according to the representational similarity analysis, or vice versa. For behavioral metrics that aim to characterize the trial-by-trial consistency, one challenge has been to decouple task performance and trial-by-trial consistency. DVC provides a principled way to do so. For future work, it would be interesting to systematically compare and theoretically relate DVC to other proposed similarity measures [15, 5, 18, 17]. We show that DVC reveals surprising negative correlations between (i) the classification performance of the deep networks trained on ImageNet-1k and (ii) the consistency with the neural representation in V4/IT. Also surprisingly, training the deep network adversarially or using rich datasets seem to evoke a decrease, rather than an increase of DVCs. While it is unclear how to close the gap between the image-by-image consistency of the deep networks to that of the brain, we think the following directions might be promising: (i) training networks using datasets that better resemble the stimulus statistics that drives the evolution of the primate visual system [23]; (ii) develop training procedures that better capture the stimulus noise and internal noise of the brain [42], as well as low level properties of the visual system (e.g., optics and foveation).
Limitations First, our results are limited by the number of monkey subjects in the datasets and the number of simultaneously recorded neurons in V4 and IT. Future larger neural datasets would allow for more accurate estimates of DVCs. Second, despite various adjustments in dimensionality reduction and DV decoding that we have experimented with, there may be factors that we have not taken into account that limit the scope and applicability of the results. For example, It is possible that the high-dimensionality of the feature space of some models affected the estimation of the DV. However, the DVCs of these models with the monkeys are not systematically lower, thus it is unlikely that the they are underestimated. Third, while monkeys provide access to neural recordings, the objects shown in the experiments might not have the same behavioral relevance as they do for humans. Thus, caution should be taken when attempting to generalize the result to humans.
Appendix A Method details
A.1 Split normalization recovers true DVC
Assume the noisy DVs:
$\text{DV}_{A}$ $=s_{A}+\varepsilon_{A}\quad$ $\text{DV}_{B}$ $=s_{B}+\varepsilon_{B}$
Assume mean-centered and all signal-noise and noise-noise covariances vanish:
$\mathbb{E}[s_{A}]=\mathbb{E}[s_{B}]=\mathbb{E}[\varepsilon_{A}]=\mathbb{E}[\varepsilon_{B}]=0$
$\mathrm{Cov}(s_{A},\varepsilon_{A})=\mathrm{Cov}(s_{A},\varepsilon_{B})=\mathrm{Cov}(s_{B},\varepsilon_{A})=\mathrm{Cov}(s_{B},\varepsilon_{B})=\mathrm{Cov}(\varepsilon_{B},\varepsilon_{B})=0$
Note:
$\mathrm{Var}(s_{A})=\sigma_{A}^{2},\quad\mathrm{Var}(s_{B})=\sigma_{B}^{2},\quad\mathrm{Cov}(s_{A},s_{B})=\rho_{\text{true}}\sigma_{A}\sigma_{B}$
$\mathrm{Var}(\varepsilon_{A})=\sigma_{\varepsilon_{A}}^{2},\quad\mathrm{Var}(\varepsilon_{B})=\sigma_{\varepsilon_{B}}^{2}$
Then:
$\mathrm{Cov}(\text{DV}_{A},\text{DV}_{B})=\mathrm{Cov}(s_{A},s_{B})=\rho_{\text{true}}\sigma_{A}\sigma_{B}$
$\mathrm{Var}(\text{DV}_{A})=\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2},\quad\mathrm{Var}(\text{DV}_{B})=\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2}$
So the observed correlation is:
$\rho_{\text{obs}}=\rho_{\text{true}}\cdot\frac{\sigma_{A}\sigma_{B}}{\sqrt{(\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2})(\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2})}}$
Now split both DVs:
$\text{DV}_{A1}=s_{A}+\varepsilon_{A1},$ $\text{DV}_{A2}=s_{A}+\varepsilon_{A2},\quad$ $\text{DV}_{B1}=s_{B}+\varepsilon_{B1},$ $\text{DV}_{B2}=s_{B}+\varepsilon_{B2}$
Assuming independent, identically distributed splits, and zero-mean and zero-covariances as before:
$\mathrm{Var}(\text{DV}_{A1})=\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2},\quad\mathrm{Cov}(\text{DV}_{A1},\text{DV}_{A2})=\sigma_{A}^{2}$
So the within-observer reliability is:
$\rho(\text{DV}_{A1},\text{DV}_{A2})=\frac{\sigma_{A}^{2}}{\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2}}$
Likewise for B:
$\rho(\text{DV}_{B1},\text{DV}_{B2})=\frac{\sigma_{B}^{2}}{\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2}}$
And the cross-observer split correlations are:
$\rho(\text{DV}_{Ai},\text{DV}_{Bj})=\rho_{\text{true}}\cdot\frac{\sigma_{A}\sigma_{B}}{\sqrt{(\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2})(\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2})}}$
where $i,j=(1,2)$.
Taking the geometric mean of the cross-observer split correlation gives a better estimate of $\rho_{\text{obs}}$:
$r_{\text{cross}}=\left[\rho(\text{DV}_{A1},\text{DV}_{B1})\cdot\rho(\text{DV}_{A1},\text{DV}_{B2})\cdot\rho(\text{DV}_{A2},\text{DV}_{B1})\cdot\rho(\text{DV}_{A2},\text{DV}_{B2})\right]^{1/4}=\rho_{\text{obs}},\quad$
While the geometric mean of the within-observer split correlation gives the normalization factor:
$r_{\text{self}}=\left[\rho(\text{DV}_{A1},\text{DV}_{A2})\cdot\rho(\text{DV}_{B1},\text{DV}_{B2})\right]^{1/2}=\sqrt{\frac{\sigma_{A}^{2}}{\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2}}\cdot\frac{\sigma_{B}^{2}}{\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2}}}=\frac{\sigma_{A}\sigma_{B}}{\sqrt{(\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2})(\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2})}}$
Then the noise-corrected correlation is:
$\rho_{\text{corrected}}=\frac{r_{\text{cross}}}{r_{\text{self}}}=\frac{\rho_{\text{true}}\sigma_{A}\sigma_{B}/\sqrt{(\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2})(\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2})}}{\sigma_{A}\sigma_{B}/\sqrt{(\sigma_{A}^{2}+\sigma_{\varepsilon_{A}}^{2})(\sigma_{B}^{2}+\sigma_{\varepsilon_{B}}^{2})}}=\rho_{\text{true}}$
A.2 Simulation demonstrates the relationship between bias, accuracy and Cohen’s Kappa
In order to demonstrate that this bias in decision could influence Cohen’s Kappa, we did a simple simulation. Suppose there are 10 classes with 100 samples each. The observers output a vector corresponds to the classes. An unbiased perfect observer outputs ’DV (decision variable)’ 1 for the corresponding class and 0 for all other classes (a one-hot vector). For realism, as observers make mistakes, we simply added gaussian noise to the DV output, which results in both lower Cohen’s Kappa and lower DVC. To model a biased imperfect observer, a bias is applied after DV, which is the same for all samples in the same class (e.g. 0.1 for the first class, 0.2 for the second class) etc. Varying bias levels is achieved by scaling the bias added to the output. The final output is one-hot + noise + bias.
Here, Cohen’s Kappa is directly inflated by the shared bias between two observers. On the other hand, because the bias does not affect the underlying DVs, the pre-normalization DVC is unaffected by the addition of bias. However, DVC does become systematically lower when the DVs are dominated by noise. Therefore, Cohen’s Kappa and DVC are distinct in that the former cares about the decision criterion and the latter do not. They can be seen as complimentary in certain scenarios. The simple simulation also hints at the relationship between accuracy, bias and Cohen’s Kappa. We continue this discussion in section D, where we reiterate that Cohen’s Kappa is intimately linked to accuracy.
Appendix B Model and dataset details
B.1 Model performances and choices of the penultimate layers
Table A.1: Models Trained on ImageNet-1k
[…]
, late after line=
[…]
Table A.2: Robust Models
[…]
, late after line=
[…]
Table A.3: Data-rich Models
[…]
, late after line=
[…]
While we do not have a strict criterion on selecting which models to test, we do follow certain principles. First of all, we try to cover a diverse set of model architectures and span the range of model accuracy, which is why we included older models with mediocre performances. Secondly, we try to include models that other studies have previously examined, so it is easier to compare our study to the previous studies. We did exclude some models due to time limits. We intend to examine an even more comprehensive set of models in future work.
B.2 Licenses for Third-Party Assets
The models used in this study were sourced from RobustBench, Torchvision, and Timm (PyTorch Image Models). We use these pretrained models as a cohort to study representational similarity, without referring to their individual implementation details.
We make use of publicly available datasets and pretrained models in accordance with their respective licenses:
Brain-Score/Vision dataset (Majaj et al., 2015) were used solely for non-commercial academic research. We follow the terms of use as outlined.
Models from Torchvision are provided under the BSD 3-Clause License.
Robustbench models are used under the MIT License.
Timm models are used under the Apache License 2.0.
Appendix C Implementation and verification
C.1 General information on the DVC framework
The DVC method is computationally efficient and stable, as dimensionality reduction is applied before attempting to decode the DV using LDA. All experiments were performed on Intel(R) Core i7-14700K CPU without resorting to GPU usage. Computing DVC between a pair of models take  30 seconds on average, with the total compute rounding to  25 hours.
Figure A.1: Implementation of the DVC framework. (a) The diagram of our anlaysis pipeline. To increase the accuracy of the decision variable inferred in the regime of huge dimensionality and few samples, we first reduce the dimensionality of the neural representation (or hidden layer representations in neural networks) before applying the optimal linear classifier to infer the decision variables. (b) To correct for the under-estimation of the magnitude of the inferred decision variable due to noise, we develop a normalization procedure based on estimating the effective noise from two splits of the data. See text for more details.
C.2 Verification of the robustness of DVC results
While the guiding principle behind DVC is general and intuitive, specific implementation choices carry implications on the numerical stability and robustness to different data distributions. We thus experimented with different algorithms and hyperparameter choices and found that they do not affect the main conclusions drawn in this study. Specifically, we want to verify if the choice of PC dimensions might change the conclusions in this study. First we note that with 25 PC dimensions (for each split), all the monkey recordings and model representations achieve high binary linear seprability (Figure A.2a), and that 8-way logistic regression accuracy plateaus early on (Figure A.2b). In addition, the main result is robust with varying PC dimensions (Figure A.3a,3b).
Other choices also do not affect the results. Pearson’s correlation is easily biased by extreme values. Although the DVs appear normally distributed, we substituted Pearson’s correlation with Spearman rank correlation and found that the trend persists (Figure A.3c). In addition, we tested shrink regularization for LDA, which could enhance stability of the procedure, and find that the result is robust to the choice of LDA solver (Figure A.3d).
Figure A.2: Dimensionality reduction retains task-relevant information. a. LDA score of all monkeys (red) and networks are high. b. As dimensionality increases, decoding performance using logistic regression plateaus.
Figure A.3: The main result is robust to the following choices: a. PCA dim =10; b. PCA dim = 50; c. Correlation measures (Pearson vs. Spearman); d. LDA solvers (SVD vs. eigen+shrinkage)
C.3 Split normalization under uncommon conditions
While essential to this method, the split normalization procedure could behave in unexpected ways. For example, it could result in a DVC value larger than 1 when the internal noise is high. None of the normalized correlations between different representations reached the cap. In addition, we took the absolute value before calculating the geometric means. When there is no correlation between two representations, this would bias the normalized DVC to a small positive value. None of the representations in this study fall in this range.
Appendix D More on the effect of behavioral decoding on Cohen’s Kappa
Geirhos et al. have discussed an important caveat of applying Cohen’s Kappa in details: Cohen’s Kappa is bounded by the accuracies of the subjects. Intuitively, the accuracy difference $|p_{i}-p_{j}|$ provides an upper bound on kappa. When the accuracy difference is high, one subject is often right, while the other is often wrong, then their behaviors cannot be consistent. From the original behavioral data, it is clear that human subjects perform with high accuracy but models perform poorly (Figure A.4a). While the authors used simulations to show that under the condition that the subjects act independently, accuracy difference is not necessarily correlated with model performance, it seems that the low model-human consistency is a direct result of their accuracy difference (Figure A.4b,4c).
Figure A.4: Recapitulation of results from Geirhos ’edges’ behavioral data. a. Accuracies of human subjects (red) are generally a lot higher than the models tested. b. Cohen’s Kappa between models and human subjects are low while Cohen’s Kappa between models and between human subjects are high, consistent with the original report. c. Cohen’s Kappa bounded by accuracy difference $|p_{i}-p_{j}|$.
Further, we want to verify if the high model-model consistency reported in the original work might be inflated by the low-accuracy high-bias condition caused by the choice of decoder (directly aggregating the probabilities). To test this, we take the original stimuli provided by Geirhos et al., and calculate the Cohen’s Kappa between the models by a) taking the average of the probability of the corresponding ImageNet-1k subclasses or b) training a 5-fold cross-validated logistic regression decoder on the representations in the penultimate layer. The result shows that compared to a), b) achieves higher accuracy (Figure A.5b, 5c), exhibits less bias towards certain categories (Figure A.5d,5e) and results in significantly lower model-model consistency as measured by Cohen’s Kappa (Figure A.5a).
Figure A.5: Using logistic regression decoder results in higher accuracy, lower bias and lower Cohen’s Kappa estimate compared to mean probability decoder, estimated on the ’edges’ images. a. Cohen’s Kappa estimated using a mean probability decoder is significantly higher than that estimated by a logistic regression decoder. b. Behavioral accuracy of mean probability decoder. c. Behavioral accuracy of logistic regression decoder. d. Choice histogram of the mean probabilities decoder. e. Choice histogram of the logistic regression decoder.
While the shared behavioral bias that results from aggregating probabilities from the original ImageNet-1k classes is very interesting, it does make the human-model consistency and the model-model consistency a lot more ambiguous. Therefore we think it is more suitable to use a stronger behavioral decoder or to use DVC when applicable.
Appendix E Broader Impacts
We expect DVC to be a broadly applicable approach to study the similarity of the brain and neural network models. On the positive side, the method and the results discussed in this paper could redirect community focus away from brute-force scaling and toward more targeted investigations into task-relevant representation alignment and model-brain convergence. It could in the long term lead to models that are more brain-like, thus greatly facilitating research in fields like neuroscience, cognitive science and AI interpretability and safety. However, while DVC offers a biologically grounded lens for comparing model and brain representations, promoting alignment with biological brains might inadvertently constrain models in certain domains where brain cognition is suboptimal.